# Inverse Reinforcement Learning by Estimating Expertise of Demonstrators

### v1

@@ 4 Learning from Suboptimal Data with IRL
Up to this point, we have formulated a model for each demonstrator $i$ based on the parameter $\theta^{\star}$, which defines the true reward, and parameters $\epsilon_{i}$ and $\beta_{i}$, which characterize demonstrator suboptimalities. However, we do not have access to any of these parameters, and are instead provided with a dataset of demonstrations $\mathcal{D}$. Recall that standard IRL assumes that all demonstrations come from the optimal policy $\pi_{E}$, and follows the iterative approach described in Section 2 to find it. However, applying this approach to the heterogeneous regime will yield an averaged policy, which can lead to subpar performance if a significant portion of the provided demonstrations are suboptimal.
To refine this, we introduce IRLEED, an innovative extension of the IRL framework tailored for the suboptimal setting. Firstly, Section 4.1 details how IRLEED extends the standard IRL framework, which corresponds to the predefined feature matching setting. Following this, Section 4.2 presents insightful comparisons between our IRLEED approach and the standard IRL framework, along with the related behavior cloning approach, ILEED. Finally, Section 4.3 explains how IRLEED can be applied to the generalized IRL setting, addressing the practical considerations necessary for implementing our framework in conjunction with existing IRL methodologies.
4.1 IRLEED
Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, we can jointly estimate the unknown parameters by maximizing the likelihood defined below:
@@ 4.1 IRLEED
where $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, $\beta=\{\beta_{i}\}_{i=1}^{N}$, and the probability $P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$ is given by:
$P(\tau|\theta,\epsilon_{i},\beta_{i})=p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t}).$ (8)
Under the specified demonstrator model, this likelihood serves as a suitable loss function. Specifically, rewriting the joint likelihood equivalently only over $\theta$:
$\mathcal{L}(\theta)=\max_{\epsilon,\beta}\mathcal{L}(\theta,\epsilon,\beta),$ (9)
it is easy to show that $\mathcal{L}(\theta)$ is a proper objective function.
@@
Moreover, this formulation provides us with a convenient way to derive the gradient of likelihood function $\mathcal{L}(\theta,\epsilon,\beta)$.
@@
First, note that parameter $\beta_{i}$ can be folded inside the modified soft Bellman policy of Eq. (6) as the reward estimate: $\beta_{i}(\theta+\epsilon_{i})^{\top}f(s_{t},a_{t})$. Second, we can express the likelihood as $\mathcal{L}=\sum_{i=1}^{N}\mathcal{L}_{i}$, where $\mathcal{L}_{i}=\log\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$. Setting $u_{i}=\beta_{i}(\theta+\epsilon_{i})$, we know that $\nabla_{u_{i}}\mathcal{L}_{i}=\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ $\forall i\in\{1,\ldots,N\}$ (see (Ziebart, 2010), Theorem 6.2 and Lemma A.2). Applying the chain rule completes the proof. ∎
@@
Notably, this result has an intuitive interpretation. Recall that for each demonstrator $i$, $\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ represents the difference between the empirical feature vector given samples from their policy, $\tilde{f}_{\mathcal{D}_{i}}$, and the expected feature vector under the probabilistic model, $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$. With this in mind, Eq. (11) shows that we update $\epsilon_{i}$ to directly match feature expectation with demonstrator $\pi_{i}$, just like in standard IRL. On the other hand, Eq. (10) shows that we update $\theta$ to match a weighted average over feature expectations provided by all demonstrators, where their respective precision $\beta_{i}$ determines their contribution. Finally, Eq. (12) shows that we update this precision $\beta_{i}$ to balance the expected returns of the demonstrator, $\pi_{i}$, and our probabilistic model, $\pi_{\theta,\epsilon_{i},\beta_{i}}$, under our estimate of the perceived reward, $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$. This way, when the probabilistic model outperforms the demonstration set under the estimated reward $r_{i}$, we decrease the precision $\beta_{i}$ to lower its relative performance, and vice versa.
@@
The above formulation gives us a general algorithm that extends the IRL framework described in Section 2. Specifically, IRLEED follows an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ and estimates $\theta,\epsilon,\beta$, for each demonstrator $i\in\{1,\ldots,N\}$, alternate between inferring the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ based on Eq. (6), and optimizing the parameters $\theta,\epsilon_{i},\beta_{i}$ based on Eqs. (10), (11), (12), respectively.
@@
As aforementioned, many different techniques can be used for both the inference and optimization procedures. For example, to infer the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ given $\theta,\epsilon_{i},\beta_{i}$, one can use soft versions of the standard value iteration and Q-learning algorithms depending on whether the dynamics $\mathcal{T}$ are provided or not (Bloem & Bambos, 2014). To compute feature expectations $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ based on the inferred policy, one can directly compute the expectations by utilizing the standard dynamic programming operator when the dynamics $\mathcal{T}$ are known, or estimate the expectations by utilizing Monte Carlo simulations of the MDP when the dynamics $\mathcal{T}$ are unknown.
4.2 Comparisons to Standard IRL and ILEED
@@
We can see that this likelihood maximization problem is closely related to the maximum causal entropy IRL framework described in Eqs. (1)- (5). Specifically, IRLEED makes two modification: (1) For each $i\in\{1,\ldots,N\}$, it defines individual dual variables $\theta_{i}=\theta+\epsilon_{i}$ that correspond to feature matching constraint $\tilde{f}_{\mathcal{D}_{i}}=\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$, tying the demonstrator policies together. (2) It uses a learnable parameter $\beta_{i}$ to tune the magnitude of uncertainty related to demonstrations provided in $\mathcal{D}_{i}$. Moreover, IRLEED generalizes the standard IRL framework.
@@
IRLEED recovers the standard IRL framework when we set $\epsilon_{i}=[0]^{k}$ and $\beta_{i}=1$ as constants.
@@
Given demonstrations $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ produced by suboptimal policies $\pi_{i}$ defined according to Eq. (6), let $\theta_{IRL}$ denote the naive solution to the standard IRL problem defined by Eq. 2, which treats demonstration set $\mathcal{D}$ as one homogeneous set produced by $\pi_{E}$. The performance of IRL is bounded by:
@@
where for simplicity, we assume that all demonstrators provide the same number of trajectories $M_{i}=M$.
As mentioned in Proposition 4.1, IRLEED can recover the true reward parameter $\theta^{\star}$, which outperforms $\theta_{IRL}$ unless the demonstrations provided are optimal. Unfortunately, since $\theta^{\star}$ is not the unique solution to IRLEED, we can not make guarantees on recovering the true reward parameter $\theta^{\star}$. Nonetheless, our experimental results in Section 5 demonstrate that IRLEED improves performance over IRL frameworks in both the feature matching setting described here, as well as more complex scenarios which include high dimensional input spaces and human demonstrations. Finally, we note that when the MDP is deterministic, we can express the policy without recursion:
$\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)\propto\exp(\beta_{i}(\theta+\epsilon_{i})^{\top}f(\tau)),$ (14)
where $f(\tau)=\sum_{(s,a)\in\tau}f(s,a)$, and $\pi(\tau)$ can directly map from trajectories due to the deterministic dynamics $\mathcal{T}$. Replacing the probability $P(\tau|\theta,\epsilon_{i},\beta_{i})$ with $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)$, we can see that maximizing the likelihood defined by Eq. (7) corresponds to the supervised learning approach employed by ILEED (Beliaev et al., 2022).
@@
IRLEED recovers the ILEED framework if we assume all demonstrators $i$ have full knowledge of the true reward, $\epsilon_{i}=[0]^{k}$, and ignore the underlying dynamics of the MDP.
Hence unlike ILEED, our framework provides us with a dynamics–aware solution to the problem of learning from mixtures of suboptimal demonstrations. As we demonstrate in Section 5, IRLEED can be utilized offline to achieve superior performance compared to ILEED.
4.3 Extension to Generalized IRL
While the results in the previous sections provide insight on how to learn from suboptimal demonstrations using IRL, relying on hand-crafted features $f$ can hinder our ability to express complex behavior. In general, the true reward $r$ can belong to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. In this case, one can utilize the generalized IRL objective instead:

### v2

@@ 4 Learning from Suboptimal Data with IRL
Up to this point, we have formulated a model for each demonstrator $i$ based on the parameter $\theta^{\star}$, which defines the true reward, and parameters $\epsilon_{i}$ and $\beta_{i}$, which characterize demonstrator suboptimalities. However, we do not have access to these parameters, and are instead provided with a dataset of demonstrations $\mathcal{D}$. Recall that standard IRL assumes that all demonstrations come from the optimal policy $\pi_{E}$, and follows the iterative approach described in Section 2 to find it. However, applying this approach to the heterogeneous regime will yield an averaged policy, leading to subpar performance if a significant portion of the demonstrations are suboptimal. To refine this, we introduce IRLEED, an innovative extension of the IRL framework tailored for the suboptimal setting.
IRLEED
Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, we can jointly estimate the maximum likelihood parameters:
@@
where $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, $\beta=\{\beta_{i}\}_{i=1}^{N}$, and the probability $P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$ is given by: $P(\tau|\theta,\epsilon_{i},\beta_{i})=p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t})$. Under the specified demonstrator model, this likelihood serves as a suitable loss function. Specifically, rewriting the joint likelihood equivalently only over $\theta$: $\mathcal{L}(\theta)=\max_{\epsilon,\beta}\mathcal{L}(\theta,\epsilon,\beta)$, it is easy to show that $\mathcal{L}(\theta)$ is a proper objective function.
@@
This provides us with a convenient way to derive the gradient of likelihood function $\mathcal{L}(\theta,\epsilon,\beta)$.
@@
First, the parameter $\beta_{i}$ can be folded into the modified soft Bellman policy of Eq. (6) as the reward estimate: $\beta_{i}(\theta+\epsilon_{i})^{\top}f(s_{t},a_{t})$. Second, we can express the likelihood as $\mathcal{L}=\sum_{i=1}^{N}\mathcal{L}_{i}$, where $\mathcal{L}_{i}=\log\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$. Setting $u_{i}=\beta_{i}(\theta+\epsilon_{i})$, we know that $\nabla_{u_{i}}\mathcal{L}_{i}=\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ $\forall i\in\{1,\ldots,N\}$ (see (Ziebart 2010), Theorem 6.2, Lemma A.2). Using the chain rule completes the proof. ∎
@@
This result has an intuitive interpretation. Recall that for each demonstrator $i$, $\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ represents the difference between the empirical feature vector given samples from their policy, $\tilde{f}_{\mathcal{D}_{i}}$, and the expected feature vector under the probabilistic model, $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$. With this in mind, Eq. (9) shows that we update $\epsilon_{i}$ to match feature expectation with demonstrator $\pi_{i}$, just like standard IRL. On the other hand, Eq. (8) shows that we update $\theta$ to match a weighted average over feature expectations provided by all demonstrators, where their respective precision $\beta_{i}$ determines their contribution. Lastly, Eq. (10) shows that we update this precision $\beta_{i}$ to balance the expected returns of the demonstrator, $\pi_{i}$, and our probabilistic model, $\pi_{\theta,\epsilon_{i},\beta_{i}}$, under our estimate of the perceived reward, $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$. This way, when the probabilistic model outperforms the demonstration set under the estimated reward $r_{i}$, we decrease the precision $\beta_{i}$ to lower its relative performance, and vice versa.
@@
This gives us a general algorithm that extends the IRL framework. Specifically, IRLEED follows an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ and estimates $\theta,\epsilon,\beta$, for each demonstrator $i\in\{1,\ldots,N\}$, alternate between inferring the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ based on Eq. (6), and optimizing the parameters $\theta,\epsilon_{i},\beta_{i}$ based on Eqs. (8), (9), (10), respectively.
@@
As aforementioned, many techniques can be used for both the inference and optimization procedures. To infer the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ given $\theta,\epsilon_{i},\beta_{i}$, one can use soft versions of the standard value iteration and Q-learning algorithms, depending on whether the dynamics $\mathcal{T}$ are provided or not (Bloem and Bambos 2014). To compute feature expectations $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ based on the inferred policy, one utilize the standard dynamic programming operator when the dynamics $\mathcal{T}$ are known, or estimate the expectations by utilizing Monte Carlo simulations of the MDP when the dynamics $\mathcal{T}$ are unknown.
Comparisons to Standard IRL and ILEED
@@
This likelihood maximization problem is closely related to the maximum causal entropy IRL framework described in Eqs. (1)- (5). Specifically, IRLEED makes two modification: (1) For each $i\in\{1,\ldots,N\}$, it defines individual dual variables $\theta_{i}=\theta+\epsilon_{i}$ that correspond to feature matching constraint $\tilde{f}_{\mathcal{D}_{i}}=\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$, tying the demonstrator policies together. (2) It uses a learnable parameter $\beta_{i}$ to tune the magnitude of uncertainty related to demonstrations provided in $\mathcal{D}_{i}$. Moreover, IRLEED generalizes standard IRL.
@@
IRLEED recovers the IRL framework when we set $\epsilon_{i}=[0]^{k}$ and $\beta_{i}=1$ as constants.
@@
Given $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ produced by suboptimal policies $\pi_{i}$ according to Eq. (6), let $\theta_{IRL}$ denote the naive solution to the IRL problem defined by Eq. 2, which treats demonstration set $\mathcal{D}$ as one homogeneous set produced by $\pi_{E}$. The performance of IRL is bounded by:
@@
where we assume that all demonstrators provide the same number of trajectories $M_{i}=M$.
As mentioned in Proposition 4.1, IRLEED can recover the true reward parameter $\theta^{\star}$, which outperforms $\theta_{IRL}$ unless the demonstrations provided are optimal. Unfortunately, since $\theta^{\star}$ is not the unique solution to IRLEED, we can not make guarantees on recovering the true reward parameter $\theta^{\star}$. Nonetheless, our experimental results in Section 5 demonstrate that IRLEED improves performance over IRL frameworks in both the feature matching setting described here, as well as the general IRL setting, which can be applied to continuous control tasks with high dimensional input spaces.
Finally, we note that when the MDP is deterministic, we can express the policy without recursion: $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)\propto\exp(\beta_{i}(\theta+\epsilon_{i})^{\top}f(\tau)),$ where $f(\tau)=\sum_{(s,a)\in\tau}f(s,a)$, and $\pi(\tau)$ can directly map from trajectories due to the deterministic dynamics $\mathcal{T}$. Replacing the probability $P(\tau|\theta,\epsilon_{i},\beta_{i})$ with $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)$, maximizing the likelihood defined by Eq. (7) corresponds to the supervised learning approach employed by ILEED (Beliaev et al. 2022). Hence unlike ILEED, our framework provides a dynamics–aware solution to the problem of learning from mixtures of suboptimal demonstrations.
@@
IRLEED recovers the ILEED framework if we assume all demonstrators $i$ have knowledge of the true reward, $\epsilon_{i}=[0]^{k}$, and ignore the dynamics of the MDP.
Practical Algorithm
While the results in the previous sections provide insight on how to learn from suboptimal demonstrations using IRL, there are two main concerns: (1) Relying on hand-crafted features $f$ can hinder our ability to express complex behaviors. (2) Nothing in the loss prevents the model from overfitting by learning zero shared reward $\theta$, and a completely different reward deviation $\epsilon_{i}$, for each demonstrator $i$. To address these concerns, we can extend our approach to generalized IRL:
@@
where the reward $r$ belongs to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$, and $\psi$ is a convex regularizer that is used to prevent overfitting (Ho and Ermon 2016).
In brief, we extend Eq. (12) to the suboptimal setting by using neural networks to parameterize $r$ and $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, and implement $\ell_{2}$–regularization on the outputs of $\epsilon$ to mitigate overfitting. Our experiment results in Section 5 demonstrate that IRLEED performs well in this generalized setting, and can even be trained offline when paired with Inverse soft-Q learning (IQ), an approach that avoids the iterative process defined by learning a single Q-function, implicitly representing both reward and policy (Garg et al. 2021). We leave the details of this section to Appendix A, covering how IRLEED can be utilized alongside generalized IRL algorithms to address both of the aforementioned concerns.
Figure 2: We visualize the reward recovered by IRLEED and IRL when trained using suboptimal demonstrations. Top left shows the true reward, where the three yellow corners are terminal states. Top right shows the normalized state visitation frequency over the entire dataset. Bottom left and right show the normalized rewards recovered by IRLEED and IRL respectively. We can see that the provided demonstrations are misaligned with the ground truth reward: the state visitation frequency for the top left corner is higher due to demonstrator suboptimalities. As expected, the feature matching constraint of IRL absorbs this suboptimality. Although the reward recovered by IRL contains information about the true reward, it is incorrectly biasing the top left corner. On the other hand, IRLEED is able to remove this bias, providing a better estimate of the ground truth reward.
Figure 3: We compare the performance of the policies recovered by IRLEED and IRL. The left plot shows the relative performance of IRLEED over IRL under varying dataset settings, where the top right corner corresponds to expert data. The right plot shows the performance of both policies as we increase the accuracy $\epsilon$ of demonstrators in the dataset, corresponding to the data outlined in red in the left plot. On average, IRLEED provided a $30.3\%$ improvement over IRL.
Setting Cartpole Lunar Lander
Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED Mean IQ IRLEED ILEED IQ IRLEED
L L $0.58$ $0.97$ 1.55 $1.38$ $0.96$ 1.56 $0.46$ $0.68$ 1.28 $0.84$ $0.81$ 1.34
L H $0.70$ $0.92$ 1.00 $1.00$ $0.96$ 1.00 $0.54$ $0.84$ 1.14 $1.06$ $1.05$ 1.18
H L $0.64$ $1.03$ 1.09 $1.05$ $0.95$ 1.05 $0.57$ $0.94$ 1.30 $0.85$ $0.99$ 1.14
H H $1.00$ $0.99$ $1.00$ $0.92$ $1.00$ $1.00$ $0.86$ $0.93$ 0.98 $1.01$ $0.99$ $0.98$
Table 1: Simulated Control Tasks
We list the average return of the recovered policies (IQ, IRLEED, ILEED), relative to the best demonstrator’s return, where IQ and IRLEED are used to denote online algorithms. This is provided for $4$ dataset settings (Prec., Acc.), where we list the mean performance (Mean) of all demonstrators relative to the best. The last row (H,H) corresponds to the clean dataset. The return of the best demonstrator for each setting (top to bottom) is $260,500,411,500$ (Cartpole) and $171,213,166,277$ (Lunar Lander).
5 Experiments
In this section we evaluate how IRLEED performs when learning from suboptimal demonstrations, using experiments in both online and offline IL settings, with simulated and human-generated data. Throughout this section we compared IRLEED to maximum entropy IRL (Ziebart, Bagnell, and Dey 2010), Inverse soft-Q learning (IQ) (Garg et al. 2021), and ILEED (Beliaev et al. 2022). In total, we performed four sets of experiments: (1) Using a custom Gridworld environment, we simulated suboptimal demonstrations with varying levels of precision and accuracy, and compared the performance and recovered rewards of IRLEED and maximum entropy IRL. (2) Using the control tasks, Cartpole, and Lunar Lander, we simulated suboptimal demonstrations with varying levels of expertise and accuracy, and compared the performance of IRLEED, IQ, and ILEED in both the online and offline settings. (3) Using the Atari environments, Space Invaders, and Qbert, we combined both simulated and human demonstrations, and compared the performance of IRLEED, IQ, and ILEED in the offline setting. (4) Using the continuous control Mujoco task Hopper, we collected suboptimal demonstrations from pretrained policies, and compared the performance of IRLEED and IQ.
Implementation We utilized the codebases provided by the authors to implement ILEED and IQ, creating IRLEED ontop of the IQ algorithm by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. Since IQ learns a reward and policy directly by using a single Q function as a critic, we used our additional critic network $\epsilon_{i}$ to directly add bias to the state action function instead of the reward. For further implementation details refer to Appendix B.
1. Gridworld For the Gridworld experiments, we implemented IRL and IRLEED by utilizing stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. For IRLEED, we updated $\epsilon$ and $\beta$ only after $\theta$ converged, and repeated this for two iterations.
The reward in this setting was linearly parameterized as $r(s)=\theta^{\top}f(s)$, where $f(s)$ is a one hot encoding for each gridworld state. For each level of precision and accuracy, we generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level. We tested $121$ dataset settings, collecting $40$ trajectories from each policy, and using $100$ seeds for each setting. Note that we visualized the recovered reward $r(s,a)$ for both IRLEED and IRL under one dataset setting in Figure 2. Given demonstrations that are misaligned with the ground truth reward, we saw that the feature matching constraint of IRL absorbed this suboptimality, while IRLEED was able to remove this bias, providing us with a better estimate of the ground truth reward.
To see the effect of dataset quality on performance, we use Figure 3 to show the relative improvement of the policy recovered by IRLEED over maximum entropy IRL for each dataset setting. We can see from the left plot that IRLEED provides improvement over IRL as we decrease accuracy $\epsilon$ (by decreasing $\lambda$), especially when precision $\beta$ is high. This signifies the importance of modeling the accuracy $\epsilon$, or reward bias, in addition to the precision $\beta$, or action variance, when learning from suboptimal demonstrations. This point is further highlighted in the right plot, which shows that while both methods can achieve good performance under low precision settings, IRLEED can tolerate lower demonstrator accuracy compared to IRL.
2. Simulated Control Tasks For this experiment, each level of precision and accuracy corresponded to $3$ suboptimal policies generated according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a)))$, where $Q(s,a)$ is a pretrained critic network, $\epsilon_{i}(s,a)$ is a randomly initialized critic network, and $\alpha$ is a scaling factor which is set to zero for the high accuracy setting. For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution, whereas for the high precision setting, the stochasticity of the policy was removed by following the maximum state action value. We utilized both the online and offline setting, using $30$ seeds for each dataset setting, and displaying our results in Table 1. Note that with near binary performance of the recovered policies, the variance is due to demonstration quality, hence we report averages relative to the best demonstrator computed over all initializations.
We can see that in the online setting, IRLEED outperforms its ablated counterpart, IQ, under all $3$ suboptimal dataset settings (not including the clean dataset), for both environments. Furthermore, we see that in the more complex Lunar Lander environment, IQ can not perform as well as the best demonstrator, whereas IRLEED consistently performs on par or better than the best demonstrator. To have a fair comparison between IRLEED and its behavior cloning counterpart, ILEED, we repeat these experiments without online access to environment interactions. We can see that offline IRLEED outperforms ILEED, under all $3$ suboptimal dataset settings, for both environments. Furthermore, although ILEED has comparable performance in the simpler Cartpole environment, it performs poorly in the more complex Lunar Lander environment. Moreover, we can see that ILEED shows diminished performance when demonstrators have low accuracy $\epsilon$, signifying that it can not account for the reward biases present in these settings, as was highlighted earlier in Remark 4.5.
3. Atari with Human Data Our demonstrations comprised two unique sources, dataset A: simulated data using a pretrained expert policy and dataset B: collected data using adept human players (Kurin et al. 2017). Since the human dataset B was acquired using screenshots of web-browsers, there is a mismatch in sampling frequencies and colors compared to the simulated environment used in dataset A. This provides a challenge akin to crowd-sourcing data from varying sources, where we test if IRLEED can improve IQ when training on the combined dataset AB as opposed to dataset A alone.
The results shown in Table 2 list the return of the recovered policies for both environments (Space Invaders and Qbert), averaged over $5$ seeds. While using dataset A is enough to reach decent performance, all methods do poorly when trained exclusively on dataset B due to the mismatch between the simulation environment and the data. However, we can see that IRLEED significantly improves its performance when utilizing the combined dataset AB. Overall we see IRLEED shows a $17\%$ improvement over IQ, and a $41\%$ improvement over ILEED in the combined AB setting.
Dataset Space Invaders Qbert
ILEED IQ IRLEED ILEED IQ IRLEED
AB $802$ $755$ 910 $5511$ $8182$ 9336
A $783$ $754$ $768$ $5926$ $8092$ $8049$
B $110$ $179$ $123$ 00 $809$ $1087$
Table 2: Utilizing Demonstrations from Varying Sources
We list the average return of the recovered policies, using the simulated dataset A, human dataset B (Kurin et al. 2017), and their combination AB. We note that IRLEED maintains similar per-seed variance as IQ on AB - Space Invaders: $61$ vs $54$ (IQ), Qbert: $315$ vs $376$ (IQ).
Figure 4: We plot the mean episode return of the policies learned by IRLEED (solid) and IQ (dashed) as they train on data from a single source (right) vs a mixture source (left).
4. Mujoco For our results on the Mujoco task Hopper, we collect a single trajectory from three sources, dataset A: a pretrained expert policy (return of $3457$), dataset B: a partially trained policy (return of $2121$) and a barely trained policy C (return of $216$). Due to the complex nature of this continuous control tasks, the partially trained policies can not provide adequate demonstrations for learning. Hence our goal was to see if IRLEED could improve the capability of IQ when training on the combined dataset ABC by filtering the suboptimal demonstrations.
The results shown in Figure 4 plot the mean episode return of the policies during training, averaged over $15$ initializations. As expected, both methods do poorly when training on dataset C, while achieving improved performance when training on dataset B. However, we see IRLEED outperforms IQ when training on the combined dataset ABC since it is more robust to the suboptimal demonstrations present in the data. However, we note that there is still a gap between the desired performance we see when training on dataset A alone, which demonstrates the difficulty of the task at hand.
6 Conclusion
Summary This paper addresses the challenges of leveraging suboptimal and heterogeneous demonstrations in IL by introducing IRLEED, a novel IRL framework. By integrating a general model of demonstrator suboptimality that accounts for reward bias and action variance, with a maximum entropy IRL framework, IRLEED mitigates the limitations inherent in standard IL algorithms, which often fail to account for the diversity and imperfections in real-world data. Our key insight is that: supervised learning frameworks should account for the heterogeneity in crowd-sourced datasets by leveraging information about the source identity during training.
Limitations Although we have shown that IRLEED is a general framework that improves over IRL in a multitude of settings, design choices such as the demonstrator model and reward regularization can be further analyzed. While we do try to make our framework easy to interpret, there is inherent complexity added as we need to simultaneously learn multiple demonstrator models. We also note that apart from the Gridworld experiments, we did not analyze reward recovery as it is hard to interpret in complex environments. However, our work addresses a problem setting where achieving desirable performance is a challenge in itself, and many recovered policies are below optimal. Lastly, we have not provided guarantees that state whether IRLEED will recover a better reward estimate compared to IRL. While we believe this is a good direction for future work, our goal was to present a novel formulation easily integrated with existing IRL methods, where further theoretical analysis is only possible under additional assumptions, and not crucial given the nonconvexity of the general setting.
Acknowledgements
This research was supported by NSF ECCS grant #2419982.
Appendix A Extension to Generalized IRL
While the results in Section 4 provide insight on how to learn from suboptimal demonstrations using IRL, relying on hand-crafted features $f$ can hinder our ability to express complex behavior. In general, the true reward $r$ can belong to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. In this case, one can utilize the generalized IRL objective instead: